OpenStudy (anonymous):

formula for last 2 questions in H8P3??? can any1 help me pls??

OpenStudy (anonymous):

62 and 35 worked for me

OpenStudy (anonymous):

3RD---(-ln( (3.5-5) / (0.5-5) ) * (RPU+RON) * 3.5*10^-15) * 10^12 4TH----(-ln( (0.9-(5*1800/(RON+RPU))) / (3.5-(5*RON/(RON+RPU))) ) * ((RON|| RPU) + RON) * 3.5*10^-15) * 10^12

OpenStudy (anonymous):

2nd a, vC(0+) = Q/C,,, b, th/(0.693*C) c, t = 1.3863 *RC d, Q' = 5/12*C*vC(0+) 3rd a, VIH/VOH = exp(-t/(RC)) b, VD1max = ROFF / (ROFF+RPU) * VS c, VD1- VG2 d, VOH = VG2max + (VOL - VG2max) * exp(-t/τ)

OpenStudy (anonymous):

And can you please help me with the first question... H8P1,,,,,, I cant solve it..... Please tell me formulas with anwers if you could.

OpenStudy (anonymous):

H8P1 IMPULSE a) 0 b) 2000 c) 270.6706 d) 3030.3 e) 146.3667

OpenStudy (anonymous):

we got everything correct but last two question of h8p3 is not getting green tick.....and 62 and 35 is not working for me.? So any one post the exact answers pls.

OpenStudy (anonymous):

formulas to be used are for 2nd last problem: (VOH-VS) / (VOL-VS) ) * (RPU+RON) * CGS) * 10^12 for last problem: (-ln( (VIL-(VS*RON/(RON+RPU))) / (VOH-(VS*RON/(RON+RPU))) ) * ((RON|| RPU) + RON) * CGS) * 10^12